# Transformers as Implicit State Estimators: In-Context Learning in Dynamical Systems

### v2

@@ 3 In-Context Learning for Dynamical Systems
 $I_{\hat{X}_{\text{Curr}}}\leftarrow I_{\hat{X}_{\text{next}}}$ ;
@@ 4.1 Experimental setup
For transparency and reproducibility, we build upon the code and model released by Garg et al. [5]. We adopt curriculum learning, starting with a context length of $N=10$ and incrementing it by $2$ every $2000$ training steps until reaching $N=40$. In our setting, context length refers to the number of past measurements and associated parameters available to an algorithm at inference time. Unless stated otherwise, the presented results are obtained for the hidden state dimension set to $n=8$.
Training is performed using the Adam optimizer [12] with a learning rate of $0.0001$ and a batch size of $64$. For each training example, $x_{0}$ and the measurement matrices $H_{1},H_{2},\dots,H_{N}$ are sampled from isotropic Gaussian distributions. The process noise $q_{t}$ is sampled from $\mathcal{N}(0,Q)$, where $Q=U_{Q}\Sigma_{Q}U_{Q}^{T}$. Here, $U_{Q}$ is a randomly sampled $8\times 8$ orthonormal matrix and $\Sigma_{Q}$ is diagonal with entries drawn from the uniform distribution $\mathcal{U}[0,\sigma_{q}^{2}]$. A training curriculum gradually increases $\sigma_{q}^{2}$ over 100,000100{,}000 steps, after which it is held constant at $0.025$. Similarly, the measurement noise $r_{t}$ is sampled from $\mathcal{N}(0,R)$, where $R$ is diagonal with entries $\sigma_{1}^{2},\dots,\sigma_{m}^{2}$ drawn from $\mathcal{U}[0,\sigma_{r}^{2}]$, with $\sigma_{r}^{2}$ also increasing over the first 100,000100{,}000 training steps to a fixed value of $0.025$. Both $Q$ and $R$ are re-sampled independently for each training example.
@@ 4.2 Results on linear systems
Figure 1(a) shows the mean-squared error (MSE) between the transformer’s prediction of $x_{N}$ and the ground truth, compared against several baselines including Kalman filtering, ordinary least squares, ridge regression ($\lambda=0.05$), and stochastic gradient descent (with learning rate of $\alpha=0.01$). The rationale behind these parameter choices is detailed in Appendix E.3. The corresponding MSPD values between transformer predictions and each baseline are reported in Figure 1(b). The Kalman filter achieves the lowest MSE, as expected. As context length increases, the transformer’s performance converges to that of the Kalman filter, distancing itself away from the regression-based estimators which do not account for temporal dynamics. A notable trend observed for both ridge regression and SGD baselines is a sharp rise in MSE at small-to-intermediate context lengths, followed by recovery at longer horizons. This behavior, recurring across many experiments, is consistent with the double descent phenomenon [19] wherein models initially transition from an over-parameterized to an under-parameterized regime. When the number of input features (i.e., context length) grows beyond the model’s capacity to disambiguate noise from signal, overfitting leads to increased test error. As context continues to grow, the model eventually gains sufficient statistical leverage to suppress noise and recover performance.
@@
Our final set of experiments involving linear dynamical systems investigates the transformer’s ability to perform in-context learning when no model parameters (i.e., neither the state transition matrix nor the noise covariances) are provided in the context. To keep the setup minimal, we revert to scalar measurements. Figure 5 presents both the MSE and MSPD of the transformer relative to various baselines. As seen in the figure, when faced with the challenging task of both capturing the state dynamics and implicitly estimating unknown state transition matrix, the transformer struggles and achieves performance that most closely resembles performance of SGD. However, if this task is rendered somewhat easier by reducing dimension of the state vector from $n=8$ to $n=2$, thus consequently reducing the number of parameters needed to specify unknown transition matrix, the same transformer model succeeds in closely emulating performance of the Kalman filter that is provided the information withheld from the transformer. Remarkably, despite the absence of model-specific information, the transformer’s performance progressively approaches that of the Kalman filter as the context length increases. This behavior is reminiscent of the Dual Kalman Filter (DKF)[24], which alternates between estimating the hidden state and the unknown state transition matrix. Specifically, the transition matrix is treated as a latent variable and estimated via a secondary Kalman filter that uses a regressor constructed from prior state estimates. Appendices B and C present theoretical arguments and an implementation blueprint showing how a transformer can emulate the DKF in this setting.
@@ 4.3 Results on nonlinear systems
Finally, we fix the context length to $40$ and investigate how the transformer’s behavior in System 1 depends on model capacity, varying both the number of layers and the embedding dimension. The results, summarized in Tables 1 and 2, reveal that smaller transformers exhibit behavior more similar to that of SGD and Ridge Regression, while larger models yield outputs that increasingly align with those of EKF and the particle filter, i.e., methods that incorporate knowledge of the system’s dynamical structure. This transition underscores the role of model capacity in enabling the transformer to approximate sophisticated recursive inference algorithms. Notably, even modest increases in either depth or embedding dimension significantly improve the transformer’s alignment with classical filters, as measured by MSPD. These findings suggest that architectural scale is a key factor in unlocking in-context learning of dynamics-aware estimation procedures.
@@ 5 Conclusion
This work investigated whether transformers, trained via in-context learning on synthetic trajectories from randomly sampled dynamical systems, can implicitly learn to perform filtering without test-time gradient updates or explicit access to model equations. We provided constructive arguments showing that the Kalman filter can be expressed using operations natively supported by transformer architectures, and empirically demonstrated that, in linear systems, transformer predictions closely track those of the Kalman filter when given sufficient context. The nature of the learned algorithm was found to depend on both model scale and context length – smaller models or shorter contexts emulate linear regression or stochastic gradient descent, while larger models with longer contexts converge toward optimal filtering behavior.
Beyond linear systems, we demonstrated that transformers can generalize to nonlinear dynamical systems as well. In particular, we showed that they can match the performance of the Extended Kalman Filter and the particle filter, and in some settings even outperform them. These results suggest that transformers do not merely replicate a fixed algorithm but instead learn a flexible, data-driven inference strategy.
The robustness of transformer-based filtering to missing context, such as unobserved noise covariances or system parameters, further highlighting its potential as a general-purpose alternative to manually designed filters. Future work will explore extensions to temporally correlated noise and examine the emergence of internal representations that support in-context filtering.
@@ Appendix B A brief summary of the Dual Kalman Filter
where both the latent state $x_{t}$ and the transition matrix $F_{t}$ must be estimated. The Dual Kalman Filter (DKF), introduced by Wan & Nelson [24], addresses this by alternating between two Kalman filtering recursions – one for $x_{t}$ given an estimate of $F_{t}$, and one for $F_{t}$ given an estimate of $x_{t}$.
@@ F.1 System 1: Nonlinear State Transitions with $\tanh$ Dynamics
Prior work (e.g., Akyürek et al. [1]) has observed that the GeLU nonlinearity can approximate scalar multiplication and, with sufficient additive bias, act nearly as an identity function. Building on this observation, we construct an expression that closely approximates $\tanh(x)$ using GeLU activations as
@@ F.2 System 2: Target Tracking with Unknown Turn Rate
We evaluate the transformer’s in-context learning capabilities on a practical nonlinear system: maneuvering target tracking with an unknown turning rate [15]. The state vector is $x_{t}=[a_{t},\dot{a}_{t},b_{t},\dot{b}_{t},\omega_{t}]$, where $(a_{t},b_{t})$ is the position, $(\dot{a}_{t},\dot{b}_{t})$ is the velocity, and $\omega_{t}$ is the angular turn rate.
@@ F.4 Additional Non-linear State Transitions
    System II: $f_{\eta}(x)=F\operatorname{sigmoid}(2x)+q_{t}$, with $n=2$ and $m=2$, where $\operatorname{sigmoid}(\cdot)$ denotes the sigmoid function. All other system parameters match those of System 1.
@@ Appendix G Derivation of the RAW Operator
To make this work self-contained, we briefly summarize how the RAW operator can be constructed and used to emulate key computational primitives within transformer architectures. This derivation follows the treatment in Appendix C of Akyürek et al. [1], to which we refer interested readers for additional technical details and formal proofs.
@@ G.1 Embedding Layer Design and Positional Control
Following Akyürek et al. [1], attention is directed using position embeddings that encode keys and queries. These are structured as
@@ G.3 Parameterizing the RAW Operator
We aim to show that a transformer layer with parameters $\theta=\{W^{F},W_{1},W_{2},(W^{Q},W^{K},W^{V})_{m}\}$ can approximate the RAW operator. Following the construction in Akyürek et al. [1], this can be achieved using just two attention heads. Key and query matrices are set to implement the token access pattern $K(i)$, enabling the approximation of

### v3

@@ 6 $I_{y}\leftarrow(0,\ 2n+2i)$ ;
28
29 $I_{\hat{X}_{\text{Curr}}}\leftarrow I_{\hat{X}_{\text{next}}}$ ;
30
@@ 4.1 Experimental setup
For transparency and reproducibility, we build upon the code and model released by Garg et al. (2022). We adopt curriculum learning, starting with a context length of $N=10$ and incrementing it by $2$ every $2000$ training steps until reaching $N=40$. In our setting, context length refers to the number of past measurements and associated parameters available to an algorithm at inference time. Unless stated otherwise, the presented results are obtained for the hidden state dimension set to $n=8$.
Training is performed using the Adam optimizer (Kingma, 2014) with a learning rate of $0.0001$ and a batch size of $64$. For each training example, $x_{0}$ and the measurement matrices $H_{1},H_{2},\dots,H_{N}$ are sampled from isotropic Gaussian distributions. The process noise $q_{t}$ is sampled from $\mathcal{N}(0,Q)$, where $Q=U_{Q}\Sigma_{Q}U_{Q}^{T}$. Here, $U_{Q}$ is a randomly sampled $8\times 8$ orthonormal matrix and $\Sigma_{Q}$ is diagonal with entries drawn from the uniform distribution $\mathcal{U}[0,\sigma_{q}^{2}]$. A training curriculum gradually increases $\sigma_{q}^{2}$ over 100,000100{,}000 steps, after which it is held constant at $0.025$. Similarly, the measurement noise $r_{t}$ is sampled from $\mathcal{N}(0,R)$, where $R$ is diagonal with entries $\sigma_{1}^{2},\dots,\sigma_{m}^{2}$ drawn from $\mathcal{U}[0,\sigma_{r}^{2}]$, with $\sigma_{r}^{2}$ also increasing over the first 100,000100{,}000 training steps to a fixed value of $0.025$. Both $Q$ and $R$ are re-sampled independently for each training example.
@@ 4.2 Results on linear systems
Figure 1(a) shows the mean-squared error (MSE) between the transformer’s prediction of $x_{N}$ and the ground truth, compared against several baselines including Kalman filtering, ordinary least squares, ridge regression ($\lambda=0.05$), and stochastic gradient descent (with learning rate of $\alpha=0.01$). The rationale behind these parameter choices is detailed in Appendix E.3. The corresponding MSPD values between transformer predictions and each baseline are reported in Figure 1(b). The Kalman filter achieves the lowest MSE, as expected. As context length increases, the transformer’s performance converges to that of the Kalman filter, distancing itself away from the regression-based estimators which do not account for temporal dynamics. A notable trend observed for both ridge regression and SGD baselines is a sharp rise in MSE at small-to-intermediate context lengths, followed by recovery at longer horizons. This behavior, recurring across many experiments, is consistent with the double descent phenomenon (Schaeffer et al., 2023) wherein models initially transition from an over-parameterized to an under-parameterized regime. When the number of input features (i.e., context length) grows beyond the model’s capacity to disambiguate noise from signal, overfitting leads to increased test error. As context continues to grow, the model eventually gains sufficient statistical leverage to suppress noise and recover performance. This effect is more pronounced for OLS, while the degradation in Ridge regression is milder due to the stabilizing effect of regularization.
@@
Our final set of experiments involving linear dynamical systems investigates the transformer’s ability to perform in-context learning when no model parameters (i.e., neither the state transition matrix nor the noise covariances) are provided in the context. To keep the setup minimal, we revert to scalar measurements. Figure 5 presents both the MSE and MSPD of the transformer relative to various baselines. As seen in the figure, when faced with the challenging task of both capturing the state dynamics and implicitly estimating unknown state transition matrix, the transformer struggles and achieves performance that most closely resembles performance of SGD. However, if this task is rendered somewhat easier by reducing dimension of the state vector from $n=8$ to $n=2$, thus consequently reducing the number of parameters needed to specify unknown transition matrix, the same transformer model succeeds in closely emulating performance of the Kalman filter that is provided the information withheld from the transformer. Remarkably, despite the absence of model-specific information, the transformer’s performance progressively approaches that of the Kalman filter as the context length increases. This behavior is reminiscent of the Dual Kalman Filter (DKF)(Wan and Nelson, 1996), which alternates between estimating the hidden state and the unknown state transition matrix. Specifically, the transition matrix is treated as a latent variable and estimated via a secondary Kalman filter that uses a regressor constructed from prior state estimates. Appendices B and C present theoretical arguments and an implementation blueprint showing how a transformer can emulate the DKF in this setting.
@@ 4.3 Results on nonlinear systems
Finally, we fix the context length to $40$ and investigate how the transformer’s behavior in System 1 depends on model capacity, varying both the number of layers and the embedding dimension. The results, summarized in Tables 1 and 2, reveal that smaller transformers exhibit behavior more similar to that of SGD and Ridge Regression, while larger models yield outputs that increasingly align with those of EKF and the particle filter, i.e., methods that incorporate knowledge of the system’s dynamical structure. This transition underscores the role of model capacity in enabling the transformer to approximate sophisticated recursive inference algorithms. Notably, even modest increases in either depth or embedding dimension significantly improve the transformer’s alignment with classical filters, as measured by MSPD. These findings suggest that architectural scale is a key factor in unlocking in-context learning of dynamics-aware estimation procedures. In short, while the operations required for filtering are representable in principle, sufficient model capacity is necessary to learn them reliably from data.
@@ 5 Conclusion
This work investigated whether transformers, trained via in-context learning on synthetic trajectories from randomly sampled dynamical systems, can implicitly learn to perform filtering without test-time gradient updates or explicit access to model equations. We provided constructive arguments showing that the Kalman filter can be expressed using operations natively supported by transformer architectures, and empirically demonstrated that, in linear systems, transformer predictions closely track those of the Kalman filter when given sufficient context and model capacity. The nature of the learned algorithm was found to depend on both model scale and context length – smaller models or shorter contexts emulate linear regression or stochastic gradient descent, while larger models with longer contexts converge toward optimal filtering behavior.
Beyond linear systems, we demonstrated that transformers can generalize to nonlinear dynamical systems as well. In particular, we showed that they can match the performance of the Extended Kalman Filter and the particle filter, and in some settings even outperform them. These results suggest that transformers do not merely replicate a fixed algorithm but instead learn a flexible, data-driven inference strategy through in-context learning.
The robustness of transformer-based filtering to missing context, such as unobserved noise covariances or system parameters, further highlights its potential as a general-purpose alternative to manually designed filters. However, performance degrades under limited model capacity and/or missing inputs, emphasizing the importance of expressiveness and informative prompts. Future work will explore extensions to temporally correlated noise and examine the emergence of internal representations that support in-context filtering.
@@ Appendix B A brief summary of the Dual Kalman Filter
where both the latent state $x_{t}$ and the transition matrix $F_{t}$ must be estimated. The Dual Kalman Filter (DKF), introduced by Wan and Nelson (1996), addresses this by alternating between two Kalman filtering recursions – one for $x_{t}$ given an estimate of $F_{t}$, and one for $F_{t}$ given an estimate of $x_{t}$.
@@ E.4 Multi-step Prediction
E.5 Is the transformer utilizing the parameters of the state-space model?
In the main text, we investigated the transformer’s performance on linear systems when major components of the state-space model (specifically, $F$, $Q$, and $R$) were withheld from the context. Here, we take a step further and evaluate the case where all parameters of the state-space model, including the measurement matrices $H_{t}$, are excluded. That is, the transformer must predict $\hat{y}_{N}$ given only the past measurements $\hat{y}_{1},\hat{y}_{2},\dots,\hat{y}_{N-1}$ as context. As shown in Fig. 12, in this setting the performance deteriorates sharply, as indicated by the significant increase in the resulting MSE.
(a) Strategy 1
(b) Strategy 2
Figure 12: Performance of in-context learning for a system where the state dimension is 8, under two system generation strategies. As key parameters of the state-space model are omitted from the context, the transformer’s performance degrades steadily. The gap between ICL and the Kalman filter is more pronounced under Strategy 1, where inference is more challenging.
@@ F.1 System 1: Nonlinear State Transitions with $\tanh$ Dynamics
Prior work (e.g., Akyürek et al. (2023)) has observed that the GeLU nonlinearity can approximate scalar multiplication and, with sufficient additive bias, act nearly as an identity function. Building on this observation, we construct an expression that closely approximates $\tanh(x)$ using GeLU activations as
@@ F.2 System 2: Target Tracking with Unknown Turn Rate
We evaluate the transformer’s in-context learning capabilities on a practical nonlinear system: maneuvering target tracking with an unknown turning rate Piché et al. (2012). The state vector is $x_{t}=[a_{t},\dot{a}_{t},b_{t},\dot{b}_{t},\omega_{t}]$, where $(a_{t},b_{t})$ is the position, $(\dot{a}_{t},\dot{b}_{t})$ is the velocity, and $\omega_{t}$ is the angular turn rate.
@@ F.4 Additional Non-linear State Transitions
    System II: $f_{\eta}(x)=F\sigma(2x)+q_{t}$, with $n=2$ and $m=2$, where $\sigma(\cdot)$ denotes the sigmoid function. All other system parameters match those of System 1.
@@ Appendix G Derivation of the RAW Operator
To make this work self-contained, we briefly summarize how the RAW operator can be constructed and used to emulate key computational primitives within transformer architectures. This derivation follows the treatment in Appendix C of Akyürek et al. (2023), to which we refer interested readers for additional technical details and formal proofs.
@@ G.1 Embedding Layer Design and Positional Control
Following Akyürek et al. (2023), attention is directed using position embeddings that encode keys and queries. These are structured as
@@ G.3 Parameterizing the RAW Operator
We aim to show that a transformer layer with parameters $\theta=\{W^{F},W_{1},W_{2},(W^{Q},W^{K},W^{V})_{m}\}$ can approximate the RAW operator. Following the construction in Akyürek et al. (2023), this can be achieved using just two attention heads. Key and query matrices are set to implement the token access pattern $K(i)$, enabling the approximation of